OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 36 inches and a height of 24 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

Volume of A = (12 x 12 x 8) / 3 = 384 Volume of B = (36 x 36 x 24) / 3 = 10368 to get the how many times B bigger than A you just have to divide the vol of B from vol of A. So the answer is 10368 / 384 = 27 Or there is an easy way to find this. B's base is 3times bigger than A's base and B's height is 3 times bigger than A's height. lets take A's base as P and height as Q. and formula for vol of a square pyramid is (base^2 x height)/3 so B's vol / A's vol = [ (3P)^2 x 3Q / 3 ] / [ P^2 x Q / 3 ] = 27

OpenStudy (anonymous):

27 times*